(5) Show your work (a) Let X be the time in minutes that Johnny Slack is late for his Probability class. As- sume that X is a Can someone tell me why it's wrong?
Answer #1

3 = 0.125 52) *~ Umiforom (2,10) : 5x68) març , 2<10 :> P(x<5). fWdz=$ * dx = 5-2 - 5 = 0-125 B) P(s«x«cl«<7) = P(5<x<s, x-7